# ReMA: Learning to meta-think for LLMs with multi-agent reinforcement learning

### v2

@@ Abstract
Recent research on Reasoning of Large Language Models (LLMs) has sought to further enhance their performance by integrating meta-thinking—enabling models to monitor, evaluate, and control their reasoning processes for more adaptive and effective problem-solving. However, current single-agent work lacks a specialized design for acquiring meta-thinking, resulting in low efficacy. To address this challenge, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages Multi-Agent Reinforcement Learning (MARL) to elicit meta-thinking behaviors, encouraging LLMs to think about thinking. ReMA decouples the reasoning process into two hierarchical agents: a high-level meta-thinking agent responsible for generating strategic oversight and plans, and a low-level reasoning agent for detailed executions. Through iterative reinforcement learning with aligned objectives, these agents explore and learn collaboration, leading to improved generalization and robustness. Experimental results demonstrate that ReMA outperforms single-agent RL baselines on complex reasoning tasks, including competitive-level mathematical benchmarks and LLM-as-a-Judge benchmarks. Comprehensive ablation studies further illustrate the evolving dynamics of each distinct agent, providing valuable insights into how the meta-thinking reasoning process enhances the reasoning capabilities of LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in knowledge understanding and complex reasoning tasks (Chowdhery et al., 2023; Achiam et al., 2023; Anil et al., 2023; Dubey et al., 2024). The paradigm in developing LLM-based reasoning models is shifting from scaling training-time computation towards scaling test-time computation (Snell et al., 2024). Recent advancements, such as OpenAI-o1 (OpenAI, 2024), Deepseek R1 (DeepSeek-AI et al., 2025), and Gemini 2.0 Flash Thinking (DeepMind, 2025), have demonstrated that allowing LLMs to think before generating answers can significantly enhance performance and lead to the emergence of human-like reasoning patterns. These patterns like “Wait, hold on. Let me double-check” or “Let’s break this down.”, indicate that LLMs can develop a form of meta-thinking abilities that can generalize well to out-of-distribution (OOD) tasks (Xiang et al., 2025). Meta-thinking, also known as metacognitive skills (Flavell, 1979), is an ability traditionally considered uniquely human (Didolkar et al., 2024).
@@
Figure 1: Left: A construction-based method that fine-tunes LLMs using rejection sampling, searching among combinations of pre-defined templates. Middle: RL-from-base method learns to mix meta-thinking and detailed solution steps during training. Right: Our method ReMA separates the meta-thinking and reasoning steps in a multi-agent system, allowing the agents to explore efficiently and learn to collaborate.
To address these limitations, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages multi-agent reinforcement learning (MARL) to encourage LLMs to think about thinking. Our approach employs a multi-agent system (MAS) composed of a high-level meta-thinking agent, responsible for strategic oversight and instruction generation, and a low-level reasoning agent tasked with detailed executing processes based on provided guidance. We compare the inference process among the construction-based method, R1-like method, and ReMA in Fig 1. Since MAS distributes the exploration space of SARL into multiple agents, it enables each agent to explore more structurally and efficiently during training. Then we apply reinforcement learning iteratively on each agent with aligned reward functions. In this way, ReMA effectively balances the trade-off between generalization capability and exploration efficiency. As a result, they can learn to play the best of their role (either to meta-think or to follow instructions), at the present of the other agent. Unlike SFT primarily leads to memorization (Chu et al., 2025), these two agents dynamically adjust their parameters and finally induce effective and robust collaboration. Consequently, not only does ReMA alleviate the OOD issues, but it also stimulates more organized learning of meta-thinking due to the hierarchy design.
To our knowledge, we are the first to formally define the concept of a multi-agent meta-thinking reasoning process (MAMRP) and optimize it through (multi-agent) reinforcement learning. Furthermore, we conduct extensive experiments on both math reasoning and LLM-as-a-Judge tasks. Across three backbone pretrained models on these benchmarks, ReMA achieves the highest average performance under the same training dataset. Additionally, we conduct comprehensive ablations to illustrate the evolving dynamics of each distinct agent. For instance, we observed unexpected interaction patterns characterized by a role reversal under different reward settings. These findings can provide valuable insights into how the meta-thinking reasoning process enhances the reasoning capabilities of LLMs.
@@ 2.1 Vanilla Reasoning Process (VRP)
The autoregressive decoding process of a language model can be formalized as a series of sequential token predictions, where the overall probability is equal to the product of the probabilities at each individual step. Given a language model $\pi_{\theta}$ parameterized by $\theta$, and a task prompt $\mathbf{x}=(x_{0},x_{1},\ldots,x_{N})$ of $N$ tokens, the vanilla reasoning process (VRP) directly generates a response $\mathbf{y}=(y_{0},y_{1},\ldots,y_{L})$ of $L$ tokens. The probability of generating $\mathbf{y}$ is
$\pi_{\theta}(\mathbf{y}|\mathbf{x})=\prod\limits_{i=0}^{L}\pi_{\theta}(y_{i}|x_{0},x_{1},\ldots x_{N},y_{0},\ldots,y_{i-1}).$
In the specific case where the response contains structured intermediate reasoning steps, the process is then known as chain-of-thought (CoT) (Wei et al., 2022), and the corresponding directed graph is as follows:
@@
2.2 Training VRP: Supervised Learning and Reinforcement Learning
Recently, there has been growing interest in uncovering this kind of reasoning process through RL on strong base models such as DeepSeek-V3-Base (Liu et al., 2024) and Qwen2.5-Math-7B (Yang et al., 2024). The decoding strategy of LLM follows the vanilla reasoning process and can further be formulated as a deterministic token-level Markov Decision process (Wang et al., 2024a). The optimization loss function can be denoted as follows:
$\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\theta}}\left[-\mathcal{R}(\mathbf{y},\mathbf{y}^{*})\right].$ (2)
where $\mathcal{R}(\cdot,\cdot)$ represents a reward function comparing generated answer $\mathbf{y}$ with the golden answer $\mathbf{y^{*}}$ for any question $\mathbf{x}$ sampled from dataset $\mathcal{D}$. To compute the gradient of the loss function with respect to the policy $\pi_{\theta}$, the REINFORCE algorithm (Williams, 1992) applies the log-derivative trick. Specifically, it computes the gradient as follows:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\theta}}\left[-\nabla_{\theta}\log\pi_{\theta}(\mathbf{y}\mid\mathbf{x})\mathcal{R}(\mathbf{y},\mathbf{y}^{*})\right].$
Since strict online sampling and gradient calculation can be computationally expensive, proximal policy optimization (PPO) (Schulman et al., 2017) is widely adopted for its high training stability and sampling efficiency. The PPO loss function is:
$\mathcal{L}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\text{old}}}\left[\min\left(\rho(\mathbf{y})A^{\pi_{\text{old}}},\text{clip}(\rho(\mathbf{y}),1-\epsilon,1+\epsilon)A^{\pi_{\text{old}}}\right)\right],$ (3)
where $\rho(\mathbf{y})=\frac{\pi_{\theta}(\mathbf{y|x})}{\pi_{\text{old}}(\mathbf{y|x})}$ denotes the ratio between current and the old policy, $A^{\pi_{\text{old}}}$ is the advantage of policy $\pi_{\text{old}}$, i.e., the extra reward over the average reward. REINFORCE++ (Hu, 2025a) combines the simplicity of the REINFORCE algorithm with the enhanced training stability of PPO. While REINFORCE++ builds upon the REINFORCE framework, it further improves performance by incorporating techniques from PPO such as gradient clipping, mini-batch updates, and advantage estimation.
@@ 3 Method
In this section, we present Reinforced Meta-Thinking Agent (ReMA), a reinforcement learning method integrating meta-thinking into the reasoning process of LLM under multi-agent settings (Sec 3.1), then describe the bilateral learning process enabled by MARL (Sec 3.2), and finally discuss the significant impact that reward function design has on learning dynamics (Sec 3.3).
3.1 Deploying Meta-Thinking Reasoning Process (MRP) for LLMs
Beyond VRP mentioned in Section 2.1, recent studies (Muennighoff et al., 2025; Ye et al., 2025c) have shown that integrating meta-thinking behaviors in reasoning process can largely improve the accuracy of the final answers. Meta-thinking is broadly defined as the ability to monitor, control, and evaluate one’s cognitive processes (Flavell, 1979). Effective meta-thinking fosters self-awareness, flexible thinking, and adaptive problem-solving, enabling individuals to recognize errors, seek alternative solutions, and optimize performance. We define the Meta-thinking Reasoning Process (MRP) for both single- and multi-agents. In a single-agent setting, such a reasoning process can be formulated by
$\pi_{\theta}(\mathbf{y|x})=\pi_{\theta}(\mathbf{y|x,m})\cdot\pi_{\theta}(\mathbf{m|x}),$ (4)
where $\mathbf{m}=(m_{0},m_{1},\ldots,m_{H})$ is the output tokens of meta-thinking during solving a question, which calls LLM’s once and generates meta-thinking and following reasoning autoregressively. Such a process can be further extended to a multi-turn setting as well.
Based on this formulation, we define the meta-thinking reasoning process (MRP) below:
@@ 3.1 Deploying Meta-Thinking Reasoning Process (MRP) for LLMs
Previous works are mainly construction-based methods, i.e. sampling from predefined templates of MRP and then applying supervised fine-tuning (Yue et al., ; Xi et al., 2024; Zhao et al., 2024). These works typically fail to fully explore the exponentially large language space and are confined to predefined patterns, even when using efficient search algorithms (Yao et al., 2023; Feng et al., 2023; Snell et al., 2024). To solve these problems, we separate the process and assign different LLMs. These LLMs simulate the meta-thinking reasoning process through conversation-like interactions between two agents: one is a high-level meta-thinking agent, and the other is a low-level reasoning agent.
During a conversation, the high-level and low-level agents (namely, $\pi_{\theta_{h}}$ and $\pi_{\theta_{l}}$) act in an interleaving manner. The high-level agent serves as the meta-thinking process, summarizing and generating meta-thoughts based on the question and the history of interactions, while the low-level agent focuses on detailed problem-solving and execution, following instructions provided by its counterpart. The single-turn MRP problem within the single-agent scenario, as formally defined by Eq.4, can be extended to a multi-agent setting. We formulate the multi-agent meta-thinking reasoning process (MAMRP) as follows:
$\mathbf{y}\sim\pi_{\theta_{l}}(\mathbf{y|x,m})\cdot\pi_{\theta_{h}}(\mathbf{m|x}).$ (6)
Similarly, we present the process with a directed graph with separated agents:
@@
To illustrate this procedure, we consider a representative example: The high-level agent first conducts a meta-thinking analysis of the task’s objective, appropriate solution strategies, and problem decomposition schema. This analysis is then passed to the low-level agent as a guidance for step-wise solutions. MAMRP can also be extended to a multi-turn setting: then in subsequent turns, the high-level agent can also refine or critique the low-level agent’s outputs. The process repeats until any predefined stopping criteria are met, such as reaching the maximum number of iterations or achieving a satisfactory solution.
Moreover, as a complex reasoning system, MAMRP provides various optimization opportunities in scaling inference-time computation. We leave further discussion of these aspects in Appendix A.
@@ 3.2 Training MAMRP: A Multi-agent RL Method
Figure 2: Comparison of Training Pipelines. Left: RL training of VRP and MRP. Right: MARL training in ReMA: the high-level agent is frozen while the low-level agent is trained using generated meta-thinking, execution results, and rewards. The low-level agent is frozen while the high-level agent is trained. This cycle repeats iteratively.
Rather than supervised finetuning on predefined execution pipelines, LLM(s) must possess the flexibility to explore freely and update rapidly. Therefore it is necessary to adopt RL training. To optimize the multi-agent system proposed in Sec 3.1, we frame the problem and objective function within the RL structure. First, we define a joint hierarchical policy with sequential decisions as follows:
$\pi_{(\theta_{h},\theta_{l})}(\mathbf{y}\mid\mathbf{x})=\pi_{\theta_{l}}(\mathbf{y}\mid\mathbf{x},\mathbf{m}\sim\pi_{\theta_{h}}(\cdot\mid\mathbf{x})),$ (8)
Let $R(\mathbf{y,y}^{*})$ denote the final reward, which serves as the objective function $J(\theta_{h},\theta_{l})$ for the joint policy:
$J(\theta_{h},\theta_{l})=\mathbb{E}_{\mathbf{x},\mathbf{y}^{*}}\mathbb{E}_{\mathbf{y}\sim\pi_{(\theta_{h},\theta_{l})}}R(\mathbf{y},\mathbf{y}*).$ (9)
During optimization procedure, the high-level policy $\pi_{\theta_{h}}$ and low-level policy $\pi_{\theta_{l}}$ aim to maximize their respective rewards independently. To enhance the overall system performance, such as the final success rate $R$, their individual reward functions are designed to enhance effective collaboration, ultimately leading to improved performance. The optimization goals for agents are:

### v3

@@ Abstract
Recent research on Reasoning of Large Language Models (LLMs) has sought to further enhance their performance by integrating meta-thinking—enabling models to monitor, evaluate, and control their reasoning processes for more adaptive and effective problem-solving. However, current single-agent work lacks a specialized design for acquiring meta-thinking, resulting in low efficacy. To address this challenge, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages Multi-Agent Reinforcement Learning (MARL) to elicit meta-thinking behaviors, encouraging LLMs to think about thinking. ReMA decouples the reasoning process into two hierarchical agents: a high-level meta-thinking agent responsible for generating strategic oversight and plans, and a low-level reasoning agent for detailed executions. Through iterative reinforcement learning with aligned objectives, these agents explore and learn collaboration, leading to improved generalization and robustness. Empirical results from single-turn experiments demonstrate that ReMA outperforms single-agent RL baselines on complex reasoning tasks, including competitive-level mathematical benchmarks and LLM-as-a-Judge benchmarks. Additionally, we further extend ReMA to multi-turn interaction settings, leveraging turn-level ratio and parameter sharing to improve efficiency. Comprehensive ablation studies further illustrate the evolving dynamics of each distinct agent, providing valuable insights into how the meta-thinking reasoning process enhances the reasoning capabilities of LLMs. Our code can be found in https://github.com/ziyuwan/ReMA-public
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in knowledge understanding and complex reasoning tasks (Chowdhery et al., 2023; Achiam et al., 2023; Anil et al., 2023; Dubey et al., 2024). The paradigm in developing LLM-based reasoning models is shifting from scaling training-time computation towards scaling test-time computation (Snell et al., 2024). Recent advancements, such as OpenAI-o1 (OpenAI, 2024), Deepseek R1 (DeepSeek-AI et al., 2025), and Gemini 2.0 Flash Thinking (DeepMind, 2025), have demonstrated that allowing LLMs to think before generating answers can significantly enhance performance and lead to the emergence of human-like reasoning patterns. These patterns like “Wait, hold on.” or “Let’s break this down.”, indicate that LLMs can develop a form of meta-thinking abilities that can generalize well to out-of-distribution (OOD) tasks (Xiang et al., 2025). Meta-thinking, also known as metacognitive skills (Flavell, 1979), is an ability traditionally considered uniquely human (Didolkar et al., 2024).
@@
Figure 1: Left: A construction-based method that fine-tunes LLMs using rejection sampling, searching among combinations of pre-defined templates. Middle: R1-like method learns to mix meta-thinking and detailed reasoning steps during training. Right: Our method ReMA separates the meta-thinking and reasoning steps in a multi-agent system and updated by reinforcement learning.
To address these limitations, we introduce Reinforced Meta-thinking Agents (ReMA), a novel framework that leverages multi-agent reinforcement learning (MARL) to encourage LLMs to think about thinking. Our approach employs a multi-agent system (MAS) composed of a high-level meta-thinking agent, responsible for strategic oversight and instruction generation, and a low-level reasoning agent tasked with detailed executing processes based on provided guidance. We compare the inference process among the construction-based method, R1-like method, and ReMA in Fig. 1. Since MAS distributes the exploration space of SARL into multiple agents, it enables each agent to explore more structurally and efficiently during training. Then we apply reinforcement learning on each agent with aligned reward functions. In this way, ReMA effectively balances the trade-off between generalization capability and exploration efficiency. As a result, they can learn to play the best of their role (either to meta-think or to follow instructions), at the present of the other agent.
To our knowledge, we are the first to formally define and optimize a multi-agent meta-thinking reasoning process (MAMRP) through multi-agent reinforcement learning. Our extensive experiments span both math reasoning and LLM-as-a-Judge tasks, where ReMA consistently achieves the highest average performance across three backbone pretrained models. We further extend ReMA to multi-turn interaction settings on math reasoning tasks, implementing turn-level ratio to optimize trajectory returns and stabilize training. Through comprehensive ablation studies, we illustrate the evolving dynamics between agents, revealing unexpected interaction patterns such as role reversals under different reward settings. These findings provide valuable insights into how meta-thinking processes enhance the reasoning capabilities of LLMs.
@@ 2.1 Vanilla Reasoning Process (VRP)
The probability of generating a response $\mathbf{y}$ equals the product of its stepwise probabilities. Given a model $\pi_{\theta}$ and a prompt $\mathbf{x}=(x_{1},\dots,x_{N})$, the vanilla reasoning process (VRP) autoregressively produces a response $\mathbf{y}=(y_{1},\dots,y_{L})$ with
$\pi_{\theta}(\mathbf{y}|\mathbf{x})=\prod\limits_{l=1}^{L}\pi_{\theta}(y_{l}|x_{1},x_{2},\ldots x_{N},y_{1},\ldots,y_{l-1})=\prod\limits_{l=1}^{L}\pi_{\theta}(\mathbf{y}_{l}|\mathbf{x},\mathbf{y}_{<l})$
The response usually contains intermediate reasoning steps before arriving at the final answer, this process is also known as chain-of-thought (CoT) (Wei et al., 2022), which can be represented as:
@@
2.2 Training VRP via Reinforcement Learning
RL frames VRP decoding process as a deterministic, token‐level Markov Decision process (MDP) (Wang et al., 2024a). Its objective is
$\mathcal{J}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\mathbf{y}\sim\pi_{\theta}}\left[R(\mathbf{y},\mathbf{y}^{*})\right].$
where $R(\cdot,\cdot)$ represents a reward function comparing generated answer $\mathbf{y}$ with the golden answer $\mathbf{y^{*}}$ for any question $\mathbf{x}$ sampled from dataset $\mathcal{D}$.
To compute the gradient $\nabla_{\theta}\mathcal{J}(\theta)$, computationally efficient algorithms GRPO (Shao et al., 2024) and REINFORCE++ (Hu, 2025) are widely adopted. Take GRPO as an example, given a question-answer pair $\mathbf{x},\mathbf{y}^{*}$ and a group of $G$ generated responses $\mathbf{y}_{i}$, denote $\mathbf{y}_{i,j}$ as the $j$-th token of the $i$-th response, it optimizes the following token-level objective:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\;\{\mathbf{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid\mathbf{x})}$ (2)
[…]
where the token-level ratio $r_{i,j}(\theta)$ and the group-normalized advantage $\hat{A}_{i,j}$ are defined as:
$r_{i,j}(\theta)$ $=\frac{\pi_{\theta}\bigl{(}\mathbf{y}_{i,j}\mid\mathbf{x},\,\mathbf{y}_{i,<j}\bigr{)}}{\pi_{\theta_{\mathrm{old}}}\bigl{(}\mathbf{y}_{i,j}\mid\mathbf{x},\,\mathbf{y}_{i,<j}\bigr{)}},\hat{A}_{i,j}=\frac{R_{i}-\text{mean}(\{R_{i}\}_{i=1}^{G})}{\text{std}(\{R_{i}\}_{i=1}^{G})}.$
@@ 3 Method
In this section, we present Reinforced Meta-thinking Agents (ReMA), a RL method integrating meta-thinking into the reasoning process of LLM under multi-agent settings (Sec. 3.1), then describe the learning process enabled by MARL of single- and multi-turn LLM setting (Secs. 3.2.1 and 3.2.2).
3.1 Deploying Meta-Thinking Reasoning Process for LLMs
Beyond VRP (Sec. 2.1), recent studies (Muennighoff et al., 2025; Ye et al., 2025c) have shown that integrating meta-thinking behaviors in reasoning process can largely improve the accuracy of the final answers. By integrating Meta-thinking, ReMA decomposes problem solving into two sequential phases: a meta‑thinking phase that plans, monitors, or revises strategy, followed by a reasoning phase that produces the detailed solution. We analyse Meta-thinking Reasoning Process along two orthogonal axes—single- vs. multi‑agent and single- vs. multi‑turn.
In a single-agent setting, such a process calls LLM once and generates meta-thinking and the following reasoning autoregressively. We formulate the meta-thinking reasoning process (MRP) below:
$\mathbf{y}\sim\pi_{\theta}(\mathbf{y\mid x,m})\cdot\pi_{\theta}(\mathbf{m\mid x}),$ (3)
where $\mathbf{m}$ and $\mathbf{y}$ are the output of meta-thinking and reasoning respectively. We present the procedure as shown below:
@@ 3.1 Deploying Meta-Thinking Reasoning Process for LLMs
Exploring MRP reasoning through a single-agent approach is often inefficient, as it requires the language model to simultaneously master both meta-thinking and detailed problem-solving within one call. Prior research has demonstrated that activating different model capabilities through specialized agents significantly improves MRP exploration efficiency. To leverage this insight, we decouple meta-thinking and reasoning into two separate LLM agents: a high-level agent dedicated to generating meta-thinking, and a low-level agent focused on executing reasoning steps.
During a conversation, the high-level and low-level agents (i.e., $\pi_{h}$ and $\pi_{l}$) act in an interleaving manner. The high-level agent generates and summarizes meta-thoughts from the prompt and interaction history, while the low-level agent executes detailed problem‐solving under those instructions. We formulate the multi-agent meta-thinking reasoning process (MAMRP) as follows:
$\mathbf{y}\sim\pi_{l}(\mathbf{y}\mid\mathbf{x},\mathbf{m})\,\pi_{h}(\mathbf{m}\mid\mathbf{x}).$ (5)
While the single-turn MAMRP offers a straightforward approach, it lacks the ability to perform immediate and fine-grained cognitive switching during the reasoning process, which limits its effectiveness on complex, long-horizon planning tasks. Therefore, we extend Eq. (5) and formulate the multi-turn MAMRP as follows:
$\mathbf{y}_{T}\sim\prod_{t=1}^{T}\pi_{l}(\mathbf{y}_{t}\mid\mathbf{x},\{\mathbf{m,y}\}_{<t},\mathbf{m}_{t})\,\pi_{h}(\mathbf{m}_{t}\mid\mathbf{x},\{\mathbf{m,y}\}_{<t})$ (6)
where $T$ is the number of turns. Similarly, we present the process with a directed graph:
@@
As a complex reasoning system, MAMRP provides various optimization opportunities in scaling inference-time computation. We leave further discussion of these aspects in Appendix C.1.
@@ 3.2 Training MAMRP: A Multi-Agent RL Method
Multi‐agent RL, unlike single‐agent RL in a deterministic MDP, must contend with stochastic, non‐stationary dynamics and rewards, making optimization more challenging. We start by considering an easier case, the optimization of single-turn MAMRP.
3.2.1 Optimizing Single-turn MAMRP
To train the system from Sec. 3.1, we embed it as a Markov Game between the two agents. Suppose the two LLM agents are parameterized by $\theta_{h}$ and $\theta_{l}$, respectively. Define a joint hierarchical policy over sequential decisions $\mathbf{m}$ and $\mathbf{y}$:
$\mathbf{y}\sim\pi_{(\theta_{h},\theta_{l})}(\mathbf{y}\mid\mathbf{x}):=\pi_{\theta_{l}}(\mathbf{y}\mid\mathbf{x},\mathbf{m})\cdot\pi_{\theta_{h}}(\mathbf{m}\mid\mathbf{x}),$ (8)
Let $R(\mathbf{y,y}^{*})$ denote the final reward serves as the objective function $\mathcal{J}(\theta_{h},\theta_{l})$ for the joint policy:
$\mathcal{J}(\theta_{h},\theta_{l})=\mathbb{E}_{\mathbf{x},\mathbf{y}^{*}}\mathbb{E}_{\mathbf{y}\sim\pi_{(\theta_{h},\theta_{l})}}R(\mathbf{y},\mathbf{y}*).$ (9)
During optimization procedure, the high-level policy $\pi_{\theta_{h}}$ and low-level policy $\pi_{\theta_{l}}$ aim to maximize their respective rewards independently. The optimization goals for agents are: